OpenStudy (anonymous):

What is f(x+h) for the function f(x)=x^3+4x?

OpenStudy (anonymous):

that one is cube

hartnn (hartnn):

to get f(x+h), just replace all 'x' in f(x) by 'x+h'

OpenStudy (anonymous):

i don't understand

hartnn (hartnn):

example if f(x) = x^2+10 x f(x+h) will be (x+h)^2 + 10(x+h) note that i have replaced ALL 'x' in f(x) by x+h

hartnn (hartnn):

so, can you try replacing 'x' in f(x) by x+h for your question, x^3+4x ?

OpenStudy (anonymous):

x^3+4x+h

Parth (parthkohli):

\[f(\color{#C00}x) = \color{#C00}x^3 + 4\color{#c00}x \\ f(\color{#C00}h) = \color{#c00}h^3 + 4 \color{#C00}h \\ \vdots \]

OpenStudy (anonymous):

?????? i'm not sure

hartnn (hartnn):

wherever you 'see' x, just replace it by 'x+h' one more try , maybe ?

hartnn (hartnn):

u want 1 more example ?

OpenStudy (anonymous):

(x+h)^3+4(x+h)

hartnn (hartnn):

\(\checkmark \)

OpenStudy (anonymous):

is that right now?????

Parth (parthkohli):

That's it! Expand

hartnn (hartnn):

now, can you expand (x+h)^3 ?

hartnn (hartnn):

use \((a+b)^3 = a^3+3a^2b+3ab^2+b^3\)

Parth (parthkohli):

Just replace \(a\) by \(x\) and \(b\) by \(h\).

OpenStudy (anonymous):

x^3+2hx^2+2hx+hx^2+2h^2x+2h^2

Parth (parthkohli):

Uh, that is expanded a little too much.

OpenStudy (anonymous):

x^3+3hx^2+2h^x+2h^2+2hx

hartnn (hartnn):

where did h^3 go ?

OpenStudy (anonymous):

wheere will i get that h^3

hartnn (hartnn):

you'll have only 4 terms from that formula, right ?

Parth (parthkohli):

Four completely simplified terms.

hartnn (hartnn):

\((a+b)^3 = a^3+3a^2b+3ab^2+b^3\) from the last term b^3 of above

OpenStudy (anonymous):

i'm confused now

Parth (parthkohli):

Why?

OpenStudy (anonymous):

x^3+3x^h+3xh^2+h^3

hartnn (hartnn):

yes!

OpenStudy (anonymous):

is that right now. . .

hartnn (hartnn):

x^3+3x^2h+3xh^2+h^3

OpenStudy (anonymous):

then?

hartnn (hartnn):

so, you have (x+h)^3+4(x+h) =x^3+3x^2h+3xh^2+h^3 +4x+4h i think no further simplification possible, thats your final answer...

hartnn (hartnn):

(x+h) can be factored, but not required...

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^